6. Wakanda and Valhalla are two countries. Each have 3000 active workers who work 10 hours every day, meaning they have 30000 labor/hours in a day. The rest are pro gamers who don’t work and play video games and drink beers only. For Wakanda: Producing 1 gaming laptop requires 2 hours. Producing 1 barrel of beer requires 4 hours. For Valhalla: Producing 1 gaming laptop requires 5 hours. Producing 1 barrel of beer requires 1 hour

a) Draw the PPF of these two countries without trade.

b) Now assume Odin, king of Valhalla and T'Challa, king of Wakanda come to you and ask you to give them a trade deal that makes both countries better off. Briefly explain which country will export what to the other country and import what from the other and why? What do you suggest? You have full control over their production and consumption.

Hint: If you can increase the consumption of both countries, even by 1 unit, they will be better off and happier. Any point above the PPF curve is better, if its feasible.
Answer #1

Given,

Both Wakanda and Vahalla have 30,000 labor hours in a day

In Wakanda,

Producing 1 gaming laptop requires 2 hours. Therefore, the maximum number of gaming laptops that can be produced = 30,000/2 = 15,000 in a day

Similarly, producing 1 barrel of beer requires 4 hours. Therefore, the maximum number of barrels of beer that can be produced = 30,000/4 = 7,500 in a day

The opportunity cost of producing 1 gaming laptop = 7,500/15,000 = 0.5 barrels of beer

The opportunity cost of producing 1 barrel of beer = 15,000/7,500 = 2 gaming laptops

In Valhalla,

Producing 1 gaming laptop requires 5 hours. Therefore, the maximum number of gaming laptops that can be produced = 30,000/5 = 6,000 in a day

Similarly, producing 1 barrel of beer requires 1 hour. Therefore, the maximum number of barrels of beer that can be produced = 30,000/1 = 30,000 in a day

The opportunity cost of producing 1 gaming laptop = 30,000/6,000 = 5 barrels of beer

The opportunity cost of producing 1 barrel of beer = 6,000/30,000 = 0.2 gaming laptops

a. PPFs of Wakanda and Valhalla are as shown below

Barrels of Beer 32500 30000 27500 25000 22500 20000 17500 15000 12500 10000 Wakanda Valhalla 0 2000 4000 12000 14000 16000 60

b. From the above analysis, it can be observed that the opportunity cost of producing 1 gaming laptop is lower for Wakanda and the opportunity cost of producing 1 barrel of beer is lower for Valhalla. Therefore, Wakanda has a comparative advantage in the production of gaming laptops while Valhalla has a comparative advantage in the production of beer.

Therefore, Wakanda should specialize in the production of gaming laptops and hence export the same while importing beer.

Similarly, Valhalla should specialize in the production of beer and hence export the same while importing gaming laptops.

The terms of trade must be in between the opportunity costs of both the countries for the trade to be favourable to both the nations. Therefore, the suitable terms of exchange would be 1 gaming laptop for 1 barrel of beer

In such a case, let us assume that 10000 units of gaming laptops are exchanged for 10000 barrels of beer

After trade consumption of

a. Wakanda = 5000 gaming laptops and 10000 barrels of beer

b. Valhalla = 10000 gaming laptops and 20000 barrels of beer

32500 30000 27500 25000 22500 20000 17500 15000 12500 10000 After trade consumption of Valhalla Barrels of Beer Wakanda Valha

Through specialization and trade, both countries can consumer at a level outside their respective PPFs and hence, both the countries are better off with trade